OpenStudy (anonymous):

if 14% OF THE GROUP VOTED TO CHANGE ITS BYLAWS AND 16% VOTED AGAINST THE CHANGE, HOW MANY PEOPLE DID NOT VOTE? THE GROUP HAS 460 MEMBERS.

OpenStudy (anonymous):

This means that 30% of people voted in total Therefore 70% did not vote 70% of 460 = ( 70/100) * 460 = 7*46 =322

OpenStudy (anonymous):

THANKS